Solve the equation by completing the square. Round to the nearest hundredth if necessary.
x2-x=25
a. 25.75, –24.75
b. 5.52, –4.52
c. 4.97, 5.02
d. 5.45, –4.45

Answer by Alan3354(69443)

Solve the equation by completing the square. Round to the nearest hundredth if necessary.
x^2 - x = 25
Add c to each side.
c = (b/2a)^2 = (1/2)^2 = 1/4
=========
x^2 - x + 1/4 = 25 + 1/4 = 101/4
(x - 1/2)^2 = 101/4
---
Can you do the rest?